OpenStudy (anonymous):

can someone help me with this? Choose the correct description of the graph of the compound inequality 2x + 5 > 9 and 5x <= 30 A number line with an open circle on 2, shading to the left, and a closed circle on 6, shading to the right. A number line with an open circle on 2, a closed circle on 6, and shading in between. A number line with a closed circle on 2, shading to the left, and an open circle on 6, shading to the right. A number line with a closed circle on 2, an open circle on 6, and shading in between

OpenStudy (accessdenied):

Try solving for x in both inequalities and graphing it yourself. Choose which one sounds most accurate to how you graphed it.

OpenStudy (accessdenied):

\[ \begin{array}{rcr} \mathbf{2x + 5 > 9} & \text{and} & \mathbf{5x \le 30} \\ \newline 2x > 4 & \text{and} & x \le 6 \\ x > 2 & \text{and} & x \le 6 \\ \end{array} \] The graph looks something like this, with an open circle at 2 because it does not equal 2, and a closed circle at 6 because we do have "less than or EQUAL to" 6 We shade to the right of 2 for values greater than it, and to the right of 6 for the values smaller than 6.